# Mining-Gym: A Configurable RL Benchmarking Environment for Truck Dispatch Scheduling

### v1

@@ III-C1 Typical RL
III-C2 Decision-point based or Event-driven RL
@@
III-C3 Defining Agent-Environment interaction and the MDP
@@
    States: The state represents the current status of the mining operation. The state of the system at a given time must encode all the features needed for the agent to learn a relationship with the desired objective to be maximized. For this task the state of the system is encoded as a vector with the following components
    $s^{d}=[SA_{d},TA_{d}]$ (1)
    where,
    $SA_{d}$ represents shovel-related attributes as an encoded vector at the current decision point $d$. It includes Shovel ID, encoded in 3-bit chunks, Queue Length, representing the number of trucks waiting per shovel, and Shovel Status, a binary indicator of shovel availability (online/offline). Similarly, $TA_{d}$ captures truck-related attributes, including TruckID, encoded in 5-bit chunks, Trips complete, tracking the number of trips per truck, and Trip Status, a multi-bit representation of the truck’s current state, such as loading, transit, or maintenance (see Table: II).
@@
    Shovels $(SA_{d})$ Shovel ID 3-bit binary
    Queue Length Normalized float
    Shovel Status Binary (1 or 0)
    Trucks $(TA_{d})$ Truck ID 5-bit binary
    Trips Complete Normalized float
@@
    Truck Status Codes
@@
    TABLE II: State Space Description and Truck Status Codes
@@
    $R_{i}^{d}=r_{\text{imm}}^{d}+\big{(}r_{\text{epi}}^{d}\;\text{if}\;d=d_{T}\;\text{else}\;0\big{)}$ (2)
@@
    $r_{\text{imm}}^{d}$ $=-\alpha\cdot\hat{TT}_{\text{Avg}}-\beta\cdot\hat{Q}_{\text{Avg}_{d}}-\gamma\cdot(1-D_{\text{DivScr}})$ (3)
    $=-\alpha\cdot\frac{\sum_{j=i-k+1}^{i}w_{j}\cdot\tau_{j}}{\sum_{j=i-k+1}^{i}w_{j}}-\beta\cdot\frac{\sum_{j=i-k+1}^{i}w_{j}\cdot Q_{{SH}_{j}}}{\sum_{j=i-k+1}^{i}w_{j}}$
    $\quad-\gamma\cdot(1-D_{\text{DivScr}})$ (4)
    where $w_{j}=e^{\lambda(j-i+k)}$ represents exponential weighting, and $\alpha,\beta,\gamma$ are weighting parameters such that $\alpha+\beta+\gamma=1$.
    The components are:
    Average Trip Time:
    $\tau_{d}=\frac{1}{|\Omega(d)|}\sum_{\omega\in\Omega(d)}\mathcal{T}_{\omega}(d)$ (5)
    Shovel Queue Time:
    $Q_{{SH}_{d}}=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\frac{1}{|\Omega_{\gamma}(d)|}\sum_{\omega\in\Omega_{\gamma}(d)}Q_{\omega}(d)$ (6)
    Diversity Score:
    $D_{\text{DivScr}}=\frac{|\{\gamma\in\Gamma\mid A_{\gamma}>0\}|}{|\Gamma|}$ (7)
    where $A_{\gamma}$ is the number of trucks assigned to shovel $\gamma$ in the recent window.
    Episodic Reward Formulation: The episodic reward balances production efficiency and resource utilization:
    $r_{\text{epi}}^{d_{T}}$ $=\omega_{1}\cdot P_{\text{Efficiency}}-\omega_{2}\cdot D_{\text{Imbalance}}$ (8)
    $=\omega_{1}\cdot\min\left(1,\frac{P_{\text{Vol}}}{P_{\text{Vol}_{\text{Target}}}}\right)-\omega_{2}\cdot\frac{D_{\text{Penal}}}{\sum_{\gamma\in\Gamma}A_{\gamma}}$ (9)
    where: $P_{\text{Vol}}$ is total material produced per shift, $P_{\text{Vol}_{\text{Target}}}$ is the production target per shift. $D_{\text{Penal}}=\max_{\gamma\in\Gamma}A_{\gamma}-\min_{\gamma\in\Gamma}A_{\gamma}$ measures shovel utilization imbalance, $\omega_{1},\omega_{2}$ are weighting parameters where $\omega_{1}+\omega_{2}=1$.
    This structure balances short-term efficiency with long-term production goals, guiding RL agents in open-pit mining operations. During each episode (e.g., shifts of 12 hours), the agent takes $N_{\text{steps}}$ actions, and the discounted sum of rewards defines the return $G_{t}$:
    $G_{t}=\sum_{k=t+1}^{N_{\text{steps}}}\gamma^{k-t-1}R_{k}$
    where $\gamma$ is the discounting factor that determines the impact of future actions on the objective function. The objective is to maximize the expected return $G_{t}$ by training the agent to improve actions, ensuring trucks meet production targets and minimize queue formation:
    $\max_{\pi}\mathbb{E}[G_{t}|S_{t}=s_{t},A_{t}=a_{t}]$
    Policy: The policy defines the strategy that the agent follows to select actions (shovel IDs) based on the current state of the system. Given the state $s_{d}=[SA_{d},TA_{d}]$, which encodes the current shovel and truck attributes, the policy predicts the shovel ID to which a truck should be assigned at decision point $d$. The policy maximizes cumulative reward by optimizing truck-shovel assignments for operational efficiency, adapting over time based on rewards to meet short- and long-term production goals, ensuring efficient dispatch and shovel utilization. The policy $\pi_{\theta}(a|s)$ is a function that gives the probability of taking action $a$ (i.e., assigning a specific shovel ID) given the current state $s$:
    $\pi_{\theta}(a|s)=\mathbb{P}(A_{d}=a|S_{d}=s)$
    where $A_{d}$ is the action taken at decision point $d$, which is the shovel ID assigned to a truck, and $S_{d}=s$ is the state of the system at decision point $d$, including both shovel-related attributes $S_{A_{d}}$ and truck-related attributes $T_{A_{d}}$. This ensures that the agent’s actions evolve to maximize long-term efficiency and achieve the desired operational goals.
IV Implementation Features
@@
In this section we introduce different visualizations and other GUI-based tools that we provide with the MINING-GYM simulator, for better performance comparison and studying change in KPIs
IV-A Configuration System
@@ C. Equipment Performance and Efficiency
Equipment idle time $IDL_{E}$
@@ E. Other Parameters:
TABLE III: Overall list of variables and parameters used
Figure 3: (a) Shows a sample page from the GUI interface and the resulting configuration file (b) Shows the Real-time representative visualization of the mine-site. The screenshot displays trucks queued at shovels, with the rightmost shovel offline. It also shows trucks moving between the dumping site, crushers, and the shovels.
As discussed in the previous section, a large number of parameters must be initialized to run the simulator (see Table III and Table V). To provide an organized method for supplying this information, we have designed a GUI interface. This interface allows users to set values and distributions for stochastic parameters, which are then saved to a human-readable text file. This configuration file is used by the simulator to initialize its parameters during operation (see Fig. 3(a)).
IV-B Other Visualization Features
@@
Key Performance Indicators (KPIs) in mining process optimization include metrics like equipment utilization, cycle time, production throughput, fuel efficiency, and downtime. These metrics help evaluate the efficiency, productivity, and cost-effectiveness of mining operations, guiding improvements and strategic decisions to enhance overall performance and profitability. A key factor in KPI measurement is the time frame, which can range from shifts to daily, weekly, monthly, or yearly intervals. By default, we consider a 12-hour shift as the standard, but this can be customized as a parameter.
We include the following essential KPIs :
@@
    Total Production ($P_{VOL}$): Refers to the total quantity of material (ore + waste) excavated per shift. It is measured in tons, indicating the scale and efficiency of production activities.
    $P_{VOL}$ $=N\times L$
    where: $N$ = Total number of trips in a shift, L = Material load per trip (tons)
@@
    Trips per Hour ($\text{TPH}_{t}$): The “Trips per Hour” metric quantifies the rate at which the full haul cycle (SH-CR-DS-SH) is completed, normalized to an hourly basis. It provides a standardized measure of fleet productivity by capturing the number of new trips executed within the most recent hourly interval.
    $\text{TPH}(t)=\frac{\Delta t}{60}\left[N(t)-N(t-\Delta t)\right]$
    where $N(t)$ = Total number of trips (SH-CR-DS-SH) completed by all trucks up to time $t$, $N(t-\Delta t)$ = Total trips at $t-\Delta t$, $\Delta t$ = Measurement interval in minutes (typically 60).
    Total Number of Trips:
    $N(t)=\sum_{i=1}^{TR}C_{i}(t)$
    where $TR$ = Total number of trucks in the fleet, $C_{i}(t)$ = Trips completed by truck $i$ up to time $t$.

### v2

@@ 3.3.1 Typical RL
3.3.2 Decision-point based or Event-driven RL
@@
3.3.3 Defining Agent-Environment interaction and the MDP
@@
    States: The state represents the current status of the mining operation. The state of the system at a given time must encode all the features needed for the agent to learn a relationship with the desired objective to be maximized. For this task, the state of the system is encoded as a vector with the following components:
    $s^{d}=[SA_{d},TA_{d}],$ (1)
    where $SA_{d}$ represents shovel-related attributes as an encoded vector at the current decision point $d$. It includes Shovel ID, encoded in 3-bit chunks, Queue Length, representing the number of trucks waiting per shovel, and Shovel Status, a binary indicator of shovel availability (online/offline). Similarly, $TA_{d}$ captures truck-related attributes, including TruckID, encoded in 5-bit chunks, Trips complete, tracking the number of trips per truck, and Trip Status, a multi-bit representation of the truck’s current state, such as loading, transit, or maintenance (see Table 2).
@@
    Shovels $(SA_{d})$ Shovel ID 4-bit binary (supports up to 16 shovels)
    Queue Length Normalized float [0, 1] (bounded by $2\times N_{trucks}$)
    Shovel Status Binary (1=operational, 0=breakdown)
    Active Truck $(TA_{d,active})$ Truck ID 6-bit binary (supports up to 64 trucks)
    Trips Complete Normalized float (scaled by 500)
@@
    Fleet Context $(FC_{d})$ Fleet Avg Trips Normalized float [0, 1] (mean trips across all trucks)
    Recent Shovel Usage Vector of floats (proportion of last 10 decisions per shovel)
    Fleet Diversity Normalized float [0, 1] (Shannon entropy of decisions)
@@
    Table 2: Definition of the state-space representation used in the mining fleet simulation. Table (a) outlines the structured attributes describing the current system state, including shovel-specific data $(SA_{d})$, active truck attributes $(TA_{d,active})$, and fleet-level contextual variables $(FC_{d})$. Table (b) provides the corresponding 3-bit binary encoding scheme for truck operational statuses, used to represent each truck’s current activity or condition within the model.
@@
    $R_{i}^{d}=r_{\text{imm}}^{d}+\big(r_{\text{epi}}^{d}\;\text{if}\;d=d_{T}\;\text{else}\;0\big),$ (2)
@@
    3.4 Reward Function Formulation
    The immediate reward at decision point $d$ incorporates four penalty components:
@@ 3.4 Reward Function Formulation
    $r^{d}_{\text{imm}}=-\alpha\cdot\overline{TT}_{\text{Avg}}^{\text{norm}}-\beta\cdot\overline{Q}_{\text{Avg}}^{\text{norm}}-\gamma\cdot(1-D_{\text{DivScr}})-\delta\cdot S_{\text{streak}}$ (3)
    where $\alpha,\beta,\gamma,\delta$ are component weights. Min-max normalization is applied to the exponentially-weighted averages: $x^{\text{norm}}=\frac{x-x_{\text{min}}}{x_{\text{max}}-x_{\text{min}}}$
    $\overline{TT}_{\text{Avg}}=\sum_{j=1}^{k}\bar{w}_{j}\cdot\tau_{j},\quad\overline{Q}_{\text{Avg}}=\sum_{j=1}^{k}\bar{w}_{j}\cdot Q^{SH}_{j},\quad\bar{w}_{j}=\frac{e^{\mu\cdot j}}{\sum_{i=1}^{k}e^{\mu\cdot i}}$ (4)
    where $n=\min(k,|D|)$ is the effective window size with decay parameter $\lambda$. The components are:
    • (a)
      Average Trip Time: $\tau_{d}=\frac{1}{|\Omega(d)|}\sum_{\omega\in\Omega(d)}\mathcal{T}_{\omega}(d)$, where $\Omega(d)$ is the set of trucks completing trips and $\mathcal{T}_{\omega}(d)$ is the cycle time for truck $\omega$.
    • (b)
      Shovel Queue Time: $Q_{{SH}_{d}}=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\frac{1}{|\Omega_{\gamma}(d)|}\sum_{\omega\in\Omega_{\gamma}(d)}Q_{\omega}(d)$, where $\Gamma$ is the set of shovels and $Q_{\omega}(d)$ is the waiting time for truck $\omega$ at shovel $\gamma$.
    • (c)
      Diversity Score: $D_{\text{DivScr}}=H/H_{\max}$, where $H=-\sum_{i=1}^{|\Gamma|}p_{i}\log_{2}(p_{i}+\epsilon)$, $H_{\max}=\log_{2}(|\Gamma|)$, and $p_{i}=\text{count}_{i}/\sum_{j}\text{count}_{j}$ represents the proportion of assignments to shovel $i$ within a sliding window of size $|W_{\text{div}}|$. This Shannon entropy-based measure quantifies shovel utilization uniformity.
    • (d)
      Streak Penalty: $S_{\text{streak}}=\frac{1}{|W_{\text{streak}}|-1}\sum_{i=1}^{|W_{\text{streak}}|-1}\mathbb{1}(a_{i}=a_{i-1})$, where $W_{\text{streak}}$ represents the recent $|W_{\text{streak}}|$ decisions, $a_{i}$ is the assigned shovel at decision $i$, and $\mathbb{1}(\cdot)$ is the indicator function. This penalizes consecutive assignments to the same shovel.
    The reward function parameters used in the demonstration experiments were determined through preliminary tuning to balance multiple operational objectives: production efficiency, resource utilization uniformity, and queue management. Episodic production weight $\omega_{1}=0.4$ and diversity penalty weight $\omega_{2}=0.6$ were applied. A high–performance bonus $b_{1}=0.5$ was triggered when $P_{\text{ratio}}\geq 0.95$ and $D_{\text{final}}>0.65$, while a moderate–performance bonus $b_{2}=0.2$ was triggered when $P_{\text{ratio}}\geq 0.90$ and $D_{\text{final}}>0.50$. Immediate rewards used a $k=5$ decision–point sliding window with exponential decay $\mu=0.5$.
    Complete parameter specifications are documented in Table B.9 (Appendix B). These hyperparameter values are provided as baseline configurations; Mining-Gym users can readily adjust these parameters through configuration files to reflect domain-specific operational priorities or conduct ablation studies on individual reward components.
    The episodic reward at shift termination evaluates overall performance:
    $r_{\text{epi}}^{d_{T}}=\omega_{1}\cdot P_{\text{ratio}}-\omega_{2}\cdot(1-D_{\text{final}})+B_{\text{perf}}$ (5)
    where $P_{\text{ratio}}=\min(1,P_{\text{Vol}}/P_{\text{Vol,Target}})$ is the production ratio, $D_{\text{final}}$ is the final diversity score, and the performance bonus is:
    $B_{\text{perf}}=\begin{cases}b_{1}&\text{if }P_{\text{ratio}}\geq\theta_{1}\text{ and }D_{\text{final}}>\phi_{1}\\ b_{2}&\text{if }P_{\text{ratio}}\geq\theta_{2}\text{ and }D_{\text{final}}>\phi_{2}\\ 0&\text{otherwise}\end{cases}$ (6)
    with bonus values $b_{1}>b_{2}>0$ and thresholds $\theta_{1}>\theta_{2}$, $\phi_{1}>\phi_{2}$. The total episode return is:
    $G=\sum_{d=1}^{d_{T}-1}\gamma^{d-1}r_{\text{imm}}^{d}+\gamma^{d_{T}-1}\left(r_{\text{imm}}^{d_{T}}+r_{\text{epi}}^{d_{T}}\right)$ (7)
    where $\gamma\in[0,1]$ is the discount factor, and the RL objective maximizes expected return: $\max_{\theta}\mathbb{E}_{\tau\sim\pi_{\theta}}[G]$.
4 Implementation Features
@@
In this section, we introduce different visualizations and other GUI-based tools that we provide with the Mining-Gym simulator, for better performance comparison and studying change in KPIs.
4.1 Configuration System
@@ C. Equipment Performance and Efficiency
Equipment idle time $IDL_{E}$ -
@@ E. Other Parameters:
Table 3: List of variables and parameters used. See Appendix B for detailed parametrization and distributions.
Figure 3: (a) Mining-Gym graphical user interface (GUI) and the generated configuration file. (b) Real-time representative visualization of the mine site. The screengrab shows trucks queued at shovels, with the rightmost shovel offline, and trucks transiting between dump sites, crushers, and shovels.
As discussed in the previous section, a large number of parameters must be initialized to run the simulator (see Table 3). To provide an organized method for supplying this information, we have designed a GUI interface. This interface allows users to set values and distributions for stochastic parameters, which are then saved to a human-readable text file. This configuration file is used by the simulator to initialize its parameters during operation (see Fig. 3(a)). This file now segregates parameters into fixed infrastructural properties (sampled once per experiment, e.g., number of shovels ) and operational variabilities (stochastically sampled per shift, e.g., fuel consumption and MTBF ). Furthermore, to ensure standardized and reproducible testing across diverse challenges, failure conditions are managed through a dedicated scenario configuration file. This file defines a set of pre-defined stress tests (e.g., loss of one shovel and fifteen trucks) where specific failure parameters are set precisely, allowing for systematic and comparable evaluation of scheduler robustness.
4.2 Other Visualization Features
@@
Key Performance Indicators (KPIs): We measure mining efficiency through these essential metrics:
@@
    Total production $(\text{shift}^{-1})$: $P_{Vol}=N\times LO$, where $N$ is the total trips completed and $LO$ is ore load per trip (tons).
@@
    Trips $(\text{hour}^{-1})$ : $\mathcal{T}_{h}=\frac{60}{\Delta t}\times[N(t)-N(t-\Delta t)]$, where $N(t)$ = cumulative trips at time $t$, $N(t-\Delta t)$ = trips at time $t-\Delta t$, $\Delta t$ = measurement interval (minutes), $h$ is the hour index and $N(t)=\sum_{i=1}^{TR}C_{i}(t)$ with $TR$ = total trucks, $C_{i}(t)$ = trips by truck $i$
  3. 3.
    Max. queue length $(\text{hour}^{-1})$ : $\widehat{Q}_{SH,h}=\max_{k=1}^{K}\left(\max_{i=1}^{N_{SH}}Q_{SH,i}^{(k)}\right)$, where $h$ is a given hour, $N_{SH}$ the number of shovels, $\mathbf{Q}_{SH}^{(k)}$ the shovel-queue vector at time instant $k$ with element $Q_{SH,i}^{(k)}$ for shovel $i$ ($i=1,\dots,N_{SH}$), and $K$ the number of time instants in hour $h$. Measures bottleneck severity: a high value indicates major congestion events (i.e., long queues) occurred, which points to a potential failure point or poor real-time dispatching.
  4. 4.
    Avg. queue length $(\text{hour}^{-1})$ : $\overline{Q}_{SH,h}=\frac{1}{K}\sum_{k=1}^{K}\left(\frac{1}{N_{SH}}\sum_{i=1}^{N_{SH}}Q_{SH,i}^{(k)}\right)$. Let $h$ be a given hour, $N_{SH}$ the number of shovels, $\mathbf{Q}_{SH}^{(k)}$ the shovel-queue vector at time instance $k$ with element $Q_{SH,i}^{(k)}$ for shovel $i$ ($i=1,\dots,N_{SH}$), and $K$ the number of time-instance in hour $h$. This metric measures system utilization: a low value suggests trucks and shovels are well-utilized, while a high value suggests overall system congestion.
  5. 5.
    Cost per ton $(\text{shift}^{-1})$ : $\text{CPT}=\dfrac{C_{KW}+C_{EST}}{P_{Vol}}$
    where $C_{KW}$ = known costs, $C_{EST}$ = estimated costs
  6. 6.
    Fuel consumption : $FC=\frac{N\times F_{\text{Unit}}}{P_{\text{Vol}}}$
    where $F_{\text{Unit}}=F_{W}+F_{O}+F_{E}$ with $F_{W}$, $F_{O}$, $F_{E}$ = fuel for waste, ore, empty trips respectively
@@
5 Experimental Setup
5.0.1 Environment Design
To address scalability and computational efficiency, we configure a large-scale operational scenario with 8 shovels, 60 trucks, 5 crushers, and 3 dumping sites. This represents an upper–medium to large bauxite operation (approximately 5–10 million tonnes annual capacity). The resulting truck–to–shovel ratio of 7.5:1 is within typical industry practice (6–12 trucks per shovel), generating sufficient event density for stress-testing.
With 60 trucks completing 4–6 trips per shift at 100 t payloads, this setup supports approximately 24,000–36,000 t of daily material movement across three shifts. More importantly for benchmarking, this scale produces high event density: concurrent truck movements, simultaneous loading/dumping, and overlapping breakdown–repair cycles stress-test both the DES engine and RL decision-making under realistic operational loads.
This configuration also enables controlled scalability studies across: (i) computational performance under dense event streams, (ii) algorithmic robustness under resource degradation (e.g., losing 15–30 trucks or 2–3 shovels), and (iii) queue-management complexity under realistic contention. The scale selected represents an upper bound where RL training remains tractable (5̄00 episodes within reasonable runtimes) while still capturing the operational characteristics and challenges of industrial-scale mining logistics.
Equipment specifications are listed in Table 4, and a conceptual diagram of the minesite is presented in Fig. 4.
Equipment Model/ Make Quantity Specification (each)
Shovels CAT 6040 8 43.7 t bucket size
Trucks CAT 777 60 98.2 nominal payload
Crushers Metso Lokotrack (LT106) 5 450 t / hr capacity
Dumping site - 2 Unlimited
Table 4: Equipment List with Specifications
5.0.2 Training Setup
To ensure that the trained agent generalizes well across the testing scenarios, we not only incorporate stochastic parameters to randomize breakdown and repair times for mining equipment and trucks but also enhance the reward function by introducing queue buildup metrics (queue length) and diversity score to prevent repeatedly selecting same shovels. For training an RL policy, we have used the PPO [30] algorithm from Stable Baselines3 [27]). We have used all the default parameters and training settings as in the repository’s PPO implementation. Note the scenario parameters which help specify the failure scenarios to be simulated during training or testing. To validate Mining-Gym as a benchmarking platform and highlight the potential of RL-based truck dispatch scheduling, we present comparative results using a PPO policy trained for 500 episodes (shifts). While a proof-of-concept with limited training, these results suffice to demonstrate both the framework’s feasibility and the performance potential of learned scheduling policies versus classical heuristics.
5.0.3 Testing Setup
For the random scheduler, the dispatcher’s action $A_{d}$ is selected randomly from the available set of shovel IDs $\{1,2,\dots,SH\}$. The selection follows a uniform distribution,
$P(A_{d}=a|S_{d}=s)=\frac{1}{SH}\quad\text{for all }a\in\{1,2,\dots,SH\},$
i.e. any state $S_{d}=s$, each action $A_{d}=a$ (where $a$ is a shovel ID) is equally likely to be chosen, with a probability of $\frac{1}{SH}$.
The fixed scheduler employs a deterministic round–robin allocation strategy. Each truck $i$ is permanently assigned to a specific shovel based on its ID using a modulo operation:
$A_{d,i}=(i-1)\bmod SH$ (8)
where $i\in\{1,2,\ldots,TR\}$ represents the truck ID, and $SH$ is the total number of shovels. This creates a static, pre–determined mapping:
$P(A_{d,i}=a\mid S_{d}=s)=\begin{cases}1,&\text{if }a=(i-1)\bmod SH\\ 0,&\text{otherwise}\end{cases}$ (9)
Unlike the random scheduler, the fixed scheduler’s allocation is independent of the current system state $S_{d}$ and remains constant throughout operation, distributing trucks evenly across shovels in a cyclic pattern.
In contrast, the RL policy $\pi_{\theta}(a|s)$ is a learned policy, where the action $A_{d}=a$ is selected based on the current state $S_{d}=s$ and the parameters $\theta$. The policy is defined as:
$\pi_{\theta}(a|s)=P(A_{d}=a|S_{d}=s).$
This probability is learned through RL algorithm where the agent adjusts $\theta$ based on rewards from its actions, optimizing its decision-making process over time. We tested the schedulers under different challenging scenarios, which are governed by a dedicated scenario configuration file to ensure consistent parameter selection and repeatability across all comparative trials.
Table 5: Testing Scenarios for Scheduler Robustness Evaluation.
Scenario Failures Operational Objective
(#SH, #TR)
A (0, 0) Baseline: All equipment available (No failures).
B (0, 15) Moderate Truck Stress: 25% of the haul fleet is lost.
C (1, 0) Single Shovel Loss: Tests response to $\sim$12.5% loading capacity bottleneck.
D (1, 15) Combined Stress: Simultaneous constraints on both loading and hauling capacity.
E (2, 30) Severe Disruption: Pushes the system to critical limits (25% SH, 50% TR lost).
F (3, 0) Critical Bottleneck: Tests scheduling under severe $\sim$37.5% loading capacity loss.
These scenarios systematically vary the number and type of resource failures, replacing the manual parameter definitions. We tested the schedulers under different challenging scenarios, which are governed by a dedicated scenarios to ensure consistent parameter selection and repeatability across all comparative trials. These scenarios systematically vary the number and type of resource failures.
All failure scenarios are repeated according to a over ten repeat trials, with the inherent stochasticity in event timings due to probabilistic distributions. These repeats are then considered in the results section for statistical analysis and discussion. Complete experimental scenario parameters, including detailed equipment specifications, stochastic parameters, state space descriptions, and technical specifications for trucks, shovels, and crushers can be found in Appendix C.
Figure 5: Comparison of “Mean Trips per hour” and “Mean Truck Queue Length” over a 6-hour shift under six failure scenarios: (a) No failures $\{0,0\}$, (b) Moderate Truck Stress $\{0,15\}$, (c) Single Shovel Loss $\{1,0\}$, (d) Combined Stress $\{1,15\}$, (e) Severe Disruption $\{2,30\}$, and (f) Critical Bottleneck $\{3,0\}$. Shovel failures occurs between 100-150 mins and Truck failures occur between 150-200 mins across the scenarios. Scenarios progress from baseline operation to severe capacity constraints on loading and hauling.
6 Results and Discussion
[…]
6.1 Comparing Performance, Efficiency and behavior
The RL-PPO scheduler demonstrates clear superiority over both fixed and random approaches across all six operational scenarios (A to F), showing the efficacy of RL based dynamic scheduling solution.
6.1.1 Productivity Performance :
The RL scheduler achieves an average improvement of $5.7\%$ in total trips completed compared to the combined performance of fixed and random schedulers. This gain exhibits remarkable consistency across diverse operational conditions, ranging from $4.1\%$ in Scenario C (single shovel loss) to $8.1\%$ in Scenario D (combined shovel loss with moderate truck stress). The most substantial advantages emerge under Combined Stress (Scenario D) and Critical Bottleneck (Scenario F) conditions, demonstrating the RL agent’s ability to maximize throughput through superior real-time assignment decisions when resources are constrained.
6.1.2 Operational Efficiency :
The efficiency gains are even more pronounced than productivity improvements. The RL approach reduces the average mean maximum queue length by approximately $24.4\%$ across all scenarios. The magnitude of these reductions varies with operational constraints. The largest improvements occur in scenarios with hauling capacity limitations: Moderate Truck Stress (Scenario B) and Combined Stress (Scenario D), where RL reduces queue lengths by $45.2\%$ and $36.9\%$, respectively. These reductions highlight the RL scheduler’s strength in efficiently utilizing limited haul fleet resources and optimally balancing truck distribution. Scenarios with hard loading bottlenecks (C, E, F) exhibit more modest reductions, such as $4.4\%$ in Scenario E, as some queuing becomes unavoidable. Nevertheless, RL minimizes unnecessary waiting and significantly outperforms the random scheduler. These results demonstrate that RL provides a dual benefit: maximizing system output while improving resource efficiency, proving valuable across all operational states from baseline to severe disruptions.
6.1.3 Behavioral Characteristics of Scheduling Approaches :
The behavioral patterns of each scheduling method reveal fundamental differences in their operational dynamics. The RL scheduler’s superior performance stems from its immediate and effective response to system disturbances. At system start (Hour 0, initial queue of 52.0 trucks), RL demonstrates rapid stabilization. By Hour 1, RL reduces the average queue to $11.18$, representing a $44.8\%$ improvement compared to Fixed ($16.24$) and Random ($19.85$) schedulers. Despite substantial queue range across the six-hour period (reflecting the severity of Scenarios E and F), the RL scheduler consistently maintains the lowest mean queue length ($14.14$), indicating a stable allocation policy that adapts effectively without excessive oscillation.
The Fixed scheduler shows pronounced instability, evidenced by abrupt oscillations in queue length and a high queue range of $24.89$. Under failure scenarios C–F, simple rule logic cannot compensate for dynamic capacity loss, producing rapid queue spikes followed by sharp overcorrections. This cyclical behavior highlights a structural limitation of static scheduling in non-stationary systems. The Random scheduler performs worst overall, with the highest mean queue length ($22.43$) and large queue range ($18.11$), reflecting the absence of coordinated decision logic.coordination.
SID Performance ($\text{Mean}\pm\text{SD}$) Statistical Results (Cohen’s d)
RL Random Fixed Schedule vs. Random vs. Fixed
A 9000.00 $\pm$ 161.60 8111.11 $\pm$ 518.28 8877.56 $\pm$ 319.29 $2.392$ $0.987$
B 8622.22 $\pm$ 192.21 7900.00 $\pm$ 614.41 8555.56 $\pm$ 245.52 $1.587$ $0.21$
C 8433.33 $\pm$ 187.08 8033.11 $\pm$ 557.77 8380.89 $\pm$ 351.58 $1.015$ $0.31$
D 8566.67 $\pm$ 158.11 7622.22 $\pm$ 834.83 8377.78 $\pm$ 370.06 $1.572$ $0.664$
E 7744.44 $\pm$ 339.53 7388.89 $\pm$ 527.84 7677.78 $\pm$ 164.15 $0.801$ $0.450$
F 7822.22 $\pm$ 156.35 7622.22 $\pm$ 523.87 7766.67 $\pm$ 180.28 $0.517$ $0.329$
Table 6: Mean $\pm$ SD performance of algorithms. Cohen’s d effect size is reported for RL compared with Random and Fixed schedules. Positive d indicates RL superiority; $|d|<0.2$ indicates comparable performance.
6.2 Comparing Throughput
[…]
7 Conclusion and Future Work
Mining-Gym addresses a key gap in mining process optimization by offering a configurable, open-source benchmarking environment for evaluating reinforcement learning (RL) algorithms in truck dispatch scheduling. It combines high-fidelity discrete-event simulation (DES) modeling, seamless integration with OpenAI Gym, and comprehensive visualization tools. Empirical evaluation across six operational scenarios demonstrates that the RL-PPO scheduler achieves an average improvement of 5.7% in total trips completed compared to fixed and random scheduling baselines, with gains ranging from 4.1% under single-asset failure to 8.1% under combined stress conditions. The RL approach delivers substantial efficiency improvements, reducing mean maximum queue length by approximately 24.4% on average, with reductions reaching 45.2% . These productivity and efficiency gains translate to consistently superior throughput performance, with the RL scheduler demonstrating rapid stabilization and adaptive resource allocation even under severe multi-asset disruptions. Mining-Gym promotes reproducible research and supports RL adoption in industrial mining applications.
Future development will focus on several key areas to enhance Mining-Gym’s capabilities and research utility. Enhanced Fidelity will incorporate traffic modeling, terrain features, heterogeneous fleets, and operational constraints to improve simulation realism. Scenario Expansion will add test cases that reflect challenges such as weather variability, ore heterogeneity, and dynamic production targets. Multi-objective Benchmarks will enable evaluation of trade-offs among production, cost, equipment longevity, and environmental impact through configurable objective weighting and standardized multi-criteria evaluation protocols. Robustness Evaluation Framework will implement standardized uncertainty quantification protocols including systematic parameter perturbation testing, Monte Carlo robustness analysis, and statistical confidence interval reporting to support algorithmic research in risk-sensitive decision-making. Finally, Digital Twin Integration will connect simulations with real-time operational data for data-driven optimization and practical deployment.
Acknowledgments
Appendix A Modeling Assumptions and Abstractions
This section presents the different assumption and abstractions used to simplify the modeling process, in the current version of the simulator
  1. 1.
    Traffic setting and routing: The current version of the simulation does not incorporate complex vehicle-traffic scenarios. In the future iterations we will aim to integrate this functionality. Currently the simulation assumes a bidirectional and perfectly routed traffic network within the DES modeling, without no traffic congestion.
  2. 2.
    Geographical Distribution and Equipment Heterogeneity: Shovels are distributed across three geographical zones, with zone-specific travel times to crushers and dumps modeled through location cluster assignments. Additionally, shovels have performance classes that determine loading efficiency, affecting cycle times. These characteristics are not observable to the RL agent but influence the underlying system dynamics (travel times and loading times), creating a partially observable environment where the agent must learn optimal dispatching without directly knowing individual shovel capabilities or locations.